The students of a Vulyalaya were asked to participat.e in a competition for making and decorating penholders in the shape of a cylinder with a base, using card board. Each penholder was to be of radius 3 cm and height 10.5 cm. The Vidyalaya was to suppl y the competitors with card board. If there were 35 competitors, how much cardboard was required to be bought for the competition?

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
lbtal cardboard required = 36 x [2 x x 3x l0.5+ x (3>2] cm2 = 35(198 + 28.29) cm2 = 7920 cm2•